if Sam were twice as old as he is.he would be 40 years older than Jim. if Jim is 10 years younger than Sam, how ols is Sam?
Answer by ankor@dixie-net.com(12695)

if Sam were twice as old as he is.he would be 40 years older than Jim.
if Jim is 10 years younger than Sam, how old is Sam?
:
Let s = Sam's age now; Let j = Jim's age now
Write an equation for each statement:
:
"if Sam were twice as old as he is, he would be 40 years older than Jim."
2s = j + 40
:
"if Jim is 10 years younger than Sam,"
j = s - 10
:
how old is Sam?
Substitute (s-10) for j in the 1st equation:
2s = (s-10) + 40
:
2s - s = 30
:
s = 30 yrs old is Sam's age
;
;
Check solution in the statement (j = 20):
"if Sam were twice as old as he is.he would be 40 years older than Jim. "
2(30) = 40 + 20